# ParticleFormer: A 3D Point Cloud World Model for Multi-Object, Multi-Material Robotic Manipulation

### v1

@@ Abstract
3D world models (i.e., learning-based 3D dynamics models) offer a promising approach to generalizable robotic manipulation by capturing the underlying physics of environment evolution conditioned on robot actions. However, existing 3D world models are primarily limited to single-material dynamics using a particle-based Graph Neural Network model, and often require time-consuming 3D scene reconstruction to obtain 3D particle tracks for training. In this work, we present ParticleFormer, a Transformer-based point cloud world model trained with a hybrid point cloud reconstruction loss, supervising both global and local dynamics features in multi-material, multi-object robot interactions. ParticleFormer captures fine-grained multi-object interactions between rigid, deformable, and flexible materials, trained directly from real-world robot perception data without an elaborate scene reconstruction. We demonstrate the model’s effectiveness both in 3D scene forecasting tasks, and in downstream manipulation tasks using a Model Predictive Control (MPC) policy. In addition, we extend existing dynamics learning benchmarks to include diverse multi-material, multi-object interaction scenarios. We validate our method on six simulation and three real-world experiments, where it consistently outperforms leading baselines by achieving superior dynamics prediction accuracy and less rollout error in downstream visuomotor tasks. Experimental videos are available at https://particleformer.github.io/.
@@ 1 Introduction
These world models have a long history in robotics and control [18, 19]. Such models have been proposed with 2D (pixel space) or 3D (metric space) representations. For 3D world models, researchers have shown that by parameterizing objects as compositions of basic particles in 3D space and applying a graph-based neural dynamics (GBND) framework to learn the interaction laws among these particles, the model can effectively represent scene dynamics for downstream planning and control [20, 21, 22, 23, 24]. However, we observe that these methods constrain particle interaction learning to the topology defined by the graph, making the model sensitive to hyperparameters such as the maximum distance for establishing particle adjacency and the maximum number of neighbors. This sensitivity becomes particularly problematic in environments involving multi-material, multi-object interactions, which feature heterogeneous and complex contact dynamics. These models are also difficult to train with real-world data, often requiring time-consuming dynamic Gaussian Splatting (GSplat) [25, 21] reconstruction to obtain 3D particle tracks, which are then used to supervise training of the particle dynamics model through an MSE loss. 2D world models bypass these issues by directly learning task-agnostic representations based on images, capturing dynamics within unified compact latent embeddings directly from pixel space [11, 26, 27, 1, 28]. However, these methods typically encode the 2D appearance of the scene into a single latent vector, which can fail to capture underlying 3D environmental structure, limiting their applicability to compositional task scenes and fine-grained dynamics modeling.
@@ 4.1 Experiment Setup and Task Design
Figure 3: Qualitative Results for Dynamics Prediction. We compare one-step dynamics predictions from ParticleFormer and baseline methods. ParticleFormer demonstrates superior capability in capturing object dynamics and multi-material interactions. The rightmost column shows block-wise attention heatmaps from our method (attention weights from particle $i$(row) to particle $j$ (column)), revealing the learned interaction structures across both intra- and inter-material particles.
@@ 4.2 Qualitative Evaluation on Dynamics Prediction
We first evaluate whether ParticleFormer produces visually accurate and physically plausible dynamics predictions. Figure 3 compares one-step predictions from ParticleFormer, GBND, and Ours w/o Hybrid across three real-world tasks.
@@
In the Box Pushing task, ParticleFormer captures both translational and rotational rigid-body dynamics, while GBND produces unrealistic local distortions due to its limited receptive field, and Ours w/o Hybrid struggles to align orientation. In the Cloth Gathering task, ParticleFormer accurately models the chain of interaction from cloth lifting to granular flow, whereas GBND fails to propagate forces across the cloth, and Ours w/o Hybrid underestimates the resulting granular movement. In the Rope Sweeping task, ParticleFormer correctly predicts both rope deformation and induced granular displacement, while GBND fails to propagate motion through the rope, resulting in a split rope graph and unrealistic prediction; Ours w/o Hybrid again underpredicts the motion of indirectly affected granular media, underscoring the value of hybrid supervision.
@@ 5 Conclusion
In this paper, we present ParticleFormer, a state-of-the-art object-centric world model that achieves superior performance in challenging robotic manipulation tasks. ParticleFormer improves dynamics modeling accuracy and robustness through two key designs: a Transformer-based particle-parametrization framework that flexibly models soft interaction dependencies among environmental particles, and a hybrid point cloud supervision loss that captures both global structure and local motion details. Furthermore, we extend evaluation benchmarks from single-material to multi-material scenarios and demonstrate the effectiveness of ParticleFormer across both synthetic and real-world settings. We believe ParticleFormer provides a capable and easy-to-use dynamics modeling framework, with potential to push the boundaries of world model applications in robotic manipulation tasks.

### v4

@@ Abstract
3D world models (i.e., learning-based 3D dynamics models) offer a promising approach to generalizable robotic manipulation by capturing the underlying physics of environment evolution conditioned on robot actions. However, existing 3D world models are primarily limited to single-material dynamics using a particle-based Graph Neural Network model, and often require time-consuming 3D scene reconstruction to obtain 3D particle tracks for training. In this work, we present ParticleFormer, a Transformer-based point cloud world model trained with a hybrid point cloud reconstruction loss, supervising both global and local dynamics features in multi-material, multi-object robot interactions. ParticleFormer captures fine-grained multi-object interactions between rigid, deformable, and flexible materials, trained directly from real-world robot perception data without an elaborate scene reconstruction. We demonstrate the model’s effectiveness both in 3D scene forecasting tasks, and in downstream manipulation tasks using a Model Predictive Control (MPC) policy. In addition, we extend existing dynamics learning benchmarks to include diverse multi-material, multi-object interaction scenarios. We validate our method on six simulation and three real-world experiments, where it consistently outperforms leading baselines by achieving superior dynamics prediction accuracy and less rollout error in downstream visuomotor tasks. Experimental videos are available at particleformer.
@@ 1 Introduction
These world models have a long history in robotics and control [18, 19]. Such models have been proposed with 2D (pixel space) or 3D (metric space) representations. For 3D world models, researchers have shown that by parameterizing objects as compositions of basic particles in 3D space and applying a graph-based neural dynamics (GBND) framework to learn the interaction laws among these particles, the model can effectively represent scene dynamics for downstream planning and control [20, 21, sun2021adversarial, 22, 23, 24, sun2022plate]. However, we observe that these methods constrain particle interaction learning to the topology defined by the graph, making the model sensitive to hyperparameters such as the maximum distance for establishing particle adjacency and the maximum number of neighbors. This sensitivity becomes particularly problematic in environments involving multi-material, multi-object interactions, which feature heterogeneous and complex contact dynamics. These models are also difficult to train with real-world data, often requiring time-consuming dynamic Gaussian Splatting (GSplat) [25, 21] reconstruction to obtain 3D particle tracks, which are then used to supervise training of the particle dynamics model through an MSE loss. 2D world models bypass these issues by directly learning task-agnostic representations based on images, capturing dynamics within unified compact latent embeddings directly from pixel space [11, 26, 27, 1, 28]. However, these methods typically encode the 2D appearance of the scene into a single latent vector, which can fail to capture underlying 3D environmental structure, limiting their applicability to compositional task scenes and fine-grained dynamics modeling.
@@ 4.1 Experiment Setup and Task Design
Figure 3: Qualitative Results for Dynamics Prediction. We compare one-step dynamics predictions from ParticleFormer and baseline methods. ParticleFormer demonstrates superior capability in capturing object dynamics and multi-material interactions. The rightmost column shows block-wise attention heatmaps from our method (attention weights from particle ii(row) to particle $j$ (column)), revealing the learned interaction structures across both intra- and inter-material particles.
@@
This task involves using two robot arms to sweep scattered granular particles into a designated area with a nylon rope held at both ends. The model is only trained on scenes where the rope is surrounded by granular objects. However, during testing, we place the rope on one side and the granular objects on the other, and ask the model to perform model-based planning and control in this unseen configuration. The robots must coordinate to manipulate the shape and trajectory of the rope to surround and sweep granular material. Success in this task requires ParticleFormer to capture fine-grained rope dynamics and indirect contact-based interactions.
@@ 4.2 Qualitative Evaluation on Dynamics Prediction
We first evaluate whether ParticleFormer produces visually accurate and physically plausible dynamics predictions. Figure 3 compares one-step predictions from ParticleFormer, GBND, and Ours w/o Hybrid across three real-world tasks (blue arrows indicate the robot’s motion direction).
@@
In the Box Pushing task, ParticleFormer captures both translational and rotational dynamics, while GBND produces unrealistic local distortions due to its limited receptive field, and Ours w/o Hybrid struggles to align orientation. In the Cloth Gathering task, ParticleFormer accurately models the chain of interaction from cloth lifting to granular flow, whereas GBND fails to propagate forces across the cloth, and Ours w/o Hybrid underestimates the resulting granular movement. In the Rope Sweeping task, ParticleFormer correctly predicts both rope deformation and induced granular displacement, while GBND fails to propagate motion through the rope, resulting in a split rope graph and unrealistic prediction; Ours w/o Hybrid again underpredicts the motion of indirectly affected granular media, underscoring the value of hybrid supervision.
@@ 5 Conclusion
In this paper, we present ParticleFormer, a state-of-the-art world model that achieves superior performance in challenging robotic tasks. ParticleFormer improves dynamics modeling accuracy and robustness through two key designs: a Transformer-based particle-parametrization framework that flexibly models soft interaction dependencies among environmental particles, and a hybrid point cloud supervision loss that captures both global structure and local motion details. Furthermore, we extend evaluation benchmarks from single-material to multi-material scenarios and demonstrate the effectiveness of ParticleFormer across both synthetic and real-world settings. We believe ParticleFormer provides a capable and easy-to-use dynamics modeling framework, with potential to push the boundaries of world model applications in robotic manipulation tasks.
@@
Acknowledgment.
This work was supported in part by the U.S. National Science Foundation under Grant No. FRR-2342246.